OpenStudy (buttermequeasy):

olve the equation 2 over 3 k + 6 = −12 for k. Please show your work and solution

jhonyy9 (jhonyy9):

hence 2/(3k+6) = -12 2= -12(3k+6) 2= -36k -72 36k= -74 k= -74/36 k= -37/18

jhonyy9 (jhonyy9):

or

jhonyy9 (jhonyy9):

just (2/3k) +6 = -12 so than 2/3k= -18 so 2= - 54k and k= 2/(-54) k= - 1/27

jhonyy9 (jhonyy9):

how is correct ?

jhonyy9 (jhonyy9):

3k or 3k+6 is denominator ?